A consultant traveled in a car for 3 hours to attend a meeting. The return trip took only 2 hours (different route) because his speed was 8 miles per hour faster. The total trip was 291 miles. What was the consultant's speed each way?
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

3X+2(X+8)=291
3X+2X+16=291
5X=291-16
5X=275
X=275/5
X=55 MPH IS THE SPEED TO THE MEETING & 55+8=63 MPH.
PROOF
3*55+2*63=291
165+126=291
291=291

Answer by stanbon(75887)

A consultant traveled in a car for 3 hours to attend a meeting. The return trip took only 2 hours (different route) because his speed was 8 miles per hour faster. The total trip was 291 miles. What was the consultant's speed each way?
----------------
1st part of trip DATA;
Time = 3 hrs ; Rate = x mph ; distance = rt = 3x miles
-------------------------
Return trip DATA:
Time = 2 hrs ; Rate = (x+8) mph ; distance = rt = 2(x+8) miles
------------------------
EQUATION:
distance + distance = 291 miles
3x + 2x+16 = 291
5x = 275
x = 55 mph ( speed to the meeting)
x+8 = 63 ( speed from the meeting)
============================
Cheers,
Stan H.